OpenStudy (kmullis6):

Please help! I don't know how to solve this /: One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. : x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

OpenStudy (alekos):

First you need to solve for x

OpenStudy (kmullis6):

How do i do that with this? Sorry I have trouble with these

OpenStudy (alekos):

x - 0.6x = 384 x(1-0.6) = 384 so x = 384/(1-0.6)

OpenStudy (alekos):

now solve for x

OpenStudy (alekos):

then work out 60% of x